OpenStudy (anonymous):

An airplane travels at an average speed of 250 mph for 3 hours. How far does the airplane travel? A.83.3 mB.250 mi C.500 mi D.750 mi

OpenStudy (anonymous):

/i think its c

OpenStudy (jdoe0001):

250 miles in 1hr how long in 3hrs? well \(\large \begin{array}{ccllll} miles&hrs \\\hline\\ 250&1\\ x&3 \end{array}\implies \cfrac{250}{x}=\cfrac{1}{3}\) solve for "x"

OpenStudy (anonymous):

We know that formula Distance = Speed X time & we have given Speed = 250 & time = 3hours So Distance = 250 * 3 = 750 mi